OpenStudy (anonymous):

Solve for x: 4/3x + 2/3 = 6 A.x = 4 B.x = 3 C.x = 1 D.x = 0

hero (hero):

Combine the left side since denominators are the same

OpenStudy (anonymous):

\[\frac{4}{3}x+\frac{2}{3}=6\]multiply everything by \(3\) \[3\times (\frac{4}{3}x+\frac{2}{3})=6\times 3\] \[4x+2=18\] and now it should take two steps only subtract 2, divide by 4